# Machine learning short-ranged many-body interactions in colloidal systems using descriptors based on Voronoi cells

### v1

@@ Abstract
Machine learning (ML) strategies are opening the door to faster computer simulations, allowing us to simulate more realistic colloidal systems. Since the interactions in colloidal systems are often highly many-body, stemming from e.g. depletion and steric interactions, one of the challenges for these algorithms is capturing the many-body nature of these interactions. In this paper, we introduce a new ML-based strategy for fitting many-body interactions in colloidal systems where the many-body interaction is highly local. To this end, we develop Voronoi-based descriptors for capturing the local environment and fit the effective potential using a simple neural network. To test this algorithm, we consider a simple two-dimensional model for a colloid-polymer mixture, where the colloid-colloid interaction and colloid-polymer interactions are hard-disk like, while the polymers themselves interact as ideal gas particles. We find that a Voronoi-based description is sufficient to accurately capture the many-body nature of this system. Moreover, we find that the Pearson correlation function alone is insufficient to determine the predictive power of the network emphasizing the importance of additional metrics when assessing the quality of ML-based potentials.
@@ I Introduction
Over the last few decades, progress in colloidal particle synthesis has led to an incredible collection of colloidal building blocks ranging from a few nanometers to microns in size, and with an ever-increasing variety of shapes and complex interactions. Together with these impressive strides in colloidal synthesis, there has been an increasing desire to be able to simulate such systems accurately enough to predict their phase behavior, as well as to target their design. However, accurately modeling colloidal systems can be very computationally costly as it requires modeling the colloid core, ligands on their surface, the solvent molecules in which the colloids are dispersed, polymers in the solvent, etc. As such, even accurately simulating a small number of such particles can be computationally expensive.
One potential solution lies in the realm of machine learning (ML). Over the past decade, ML has emerged as an important tool in speeding up computer simulations of atomic and molecular systems behler2016perspective ; noe2020machine by replacing expensive density functional theory calculations on the atomic scale, with machine learned forces and/or potentials. After training, these ML-based interaction potentials are exploited in either molecular dynamics or Monte Carlo (MC) simulations, allowing researchers to access larger system sizes and longer time scales than would be accessible otherwise. For such systems, strategies are rapidly evolving and improving – from the early “standard” neural network (NN) approaches, to kernel fitting, on the fly training, and message passing neural networks behler2007generalized ; artrith2016implementation ; li2015molecular ; jacobsen2018fly ; jinnouchi2019phase ; kocer2022neural ; batatia2022mace .
@@
We find that for a single size ratio (colloid-to-polymer diameter), a single Voronoi-inspired ML model can reproduce the structure of all relevant phases (gas, liquid, crystal) over all densities. Additionally, we find that it is difficult to predict the performance of the ML models based on only the correlation between predicted and true effective interactions. In particular, high Pearson correlation coefficients ($>0.99999$) are insufficient to guarantee that the ML-based simulation will accurately reproduce the correct structure of the system. Hence, in this paper we compare the radial distribution function of the fitted model to results from simulations where the polymers are treated explicitly to assess the quality of the potential.
Figure 2: Probability to have an effective $n$-body colloidal interaction at an arbitrary point in space, for various polymer sizes $\sigma_{P}$ and various colloid and polymer packing fractions $\eta_{C}$ and $\eta_{P}$. Probability is measured using MC simulation of the $N_{C}\mu_{P}AT$ colloid-polymer ensemble. Simulations were performed for a system containing 1024 colloids.
@@ II.1 Model system
$\Phi_{\text{eff}}=\frac{1}{2}\sum_{i}^{N_{C}}\sum_{i\neq j}^{N_{C}}\phi_{CC}^{HS}(\left|\mathbf{R}_{i}-\mathbf{R}_{j}\right|)+\Phi_{\text{eff, CP}}(\left\{\mathbf{R}^{N_{C}}\right\}),$ (3)
@@
$\Phi_{\text{eff, CP}}(\left\{\mathbf{R}^{N_{C}}\right\})$ $=$ $-k_{B}Tz_{P}A_{\text{eff}}(\left\{\mathbf{R}^{N_{C}}\right\})$ (4)
$A_{\text{eff}}(\left\{\mathbf{R}^{N_{C}}\right\})$ $=$ $\int\mathrm{d}\mathbf{r}\,\exp\left[-\beta\sum_{i=1}^{N_{C}}\phi_{CP}(\left|\mathbf{R}_{i}-\mathbf{r}\right|)\right].$ (5)
@@
Here, $z_{P}$ is the polymer fugacity, $k_{B}$ the Boltzmann constant, $\beta=1/(k_{B}T)$, and $A_{\text{eff}}(\left\{\mathbf{R}^{N_{C}}\right\})$ is the effective free area available to the polymers. This effective free area can be interpreted as a weighted area of the system, where the weight of each point in space is determined by the potential energy that a polymer at that position would experience due to the colloid configuration $\{\mathbf{R}^{N_{C}}\}$. Note that Eqs. (3) and (5) are valid for any arbitrary colloid-colloid and colloid-polymer interaction, provided that there are no polymer-polymer interactions.
@@
In the case of hard-sphere colloid-polymer interactions, only points where a polymer can be placed without it overlapping with any of the colloids contribute to $A_{\text{eff, CP}}(\left\{\mathbf{R}^{N_{C}}\right\})$. As a result, $A_{\text{eff, CP}}(\left\{\mathbf{R}^{N_{C}}\right\})$ simplifies to the free area available to polymers, $A_{F}$.
@@
In this paper, we focus on two polymer sizes: $\sigma_{P}=0.4\sigma_{C}$ and $\sigma_{P}=0.8\sigma_{C}$. These polymer sizes are both far outside of the regime where the approximations of the Asakura-Oosawa pair potential hold, i.e. where the effective colloid-colloid interaction can be written purely as a two-body interaction gast1983polymer ; dijkstra1999phaseb . Hence, for both polymer sizes, $\Phi_{\text{eff, CP}}(\left\{\mathbf{R}^{N_{C}}\right\})$ incorporates at least three-body interaction contributions.
@@
The many-body nature of this potential arises due to the fact that more than two depletion zones (disks of radius $\sigma_{CP}/2$ around the colloid where polymer centers cannot enter) can overlap simultaneously, see Fig. 1. The probability $P(n)$ to find $n$ overlapping depletion zones at an arbitrary point in space depends on the colloid and polymer concentration, as well as the polymer size, as is shown in Fig. 2. The data in this figure is obtained by performing an MC simulation in the $N_{C}\mu_{P}AT$ ensemble, where polymers were treated explicitly. To ensure a constant $\mu_{P}$, we perform insertion and deletion moves of polymers via the usual Monte Carlo acceptance rule frenkel2023understanding . To compute $P(n)$, we randomly selected points within the system and determined the number of depletion zones overlapping with these points (where points located within the colloids were included in the computation as well). For each state point, we analyzed 45 different snapshots, where we considered $10^{6}$ random points per snapshot. From Fig. 2 we see that at lower polymer sizes there are mainly $1$- and $2$-body interactions, while at higher polymer sizes the system mainly shows $3$-body interactions.
@@ II.2 Capture effective potential using Voronoi cells
Our aim is to train an ML algorithm that is able to predict the effective interactions between colloids as described in Eq. (5). However, rather than predicting the free area for the entire system at once, we aim to compute the contributions to the effective potential of individual colloids. This requires rewriting $A_{\text{eff}}(\left\{\mathbf{R}^{N_{C}}\right\})$ in terms of contributions associated with separate colloids, i.e. $A_{\text{eff}}(\left\{\mathbf{R}^{N_{C}}\right\})=\sum_{i\in N_{C}}A^{i}_{\text{eff}}(\left\{\mathbf{R}^{N_{C}}\right\})$. To achieve this, each point in space must be assigned to a specific colloid, allowing the integral of Eq. (5) to be split into $N$ integrals. A natural way to tile space is by using Voronoi cells, where each point in space is assigned to the nearest colloid, allowing us to write
$A_{\text{eff}}(\left\{\mathbf{R}^{N_{C}}\right\})=\sum_{i\in N_{C}}\int_{\mathcal{V}_{i}}\mathrm{d}\mathbf{r}\,\exp\left[-\beta\sum_{i}^{N}\phi_{CP}(\mathbf{R}_{j}-\mathbf{r})\right],$
@@
$A_{\text{eff}}(\left\{\mathbf{R}^{N_{C}}\right\})=\sum_{i\in N_{C}}A^{(1)}({\mathcal{V}_{i}})+\sum_{i,j\in N_{C}}A^{(2)}({\mathcal{V}_{i}},{\mathcal{V}_{j}})+\ldots,$ (6)
@@
$A_{\text{eff}}(\left\{\mathbf{R}^{N_{C}}\right\})=\sum_{i\in N_{C}}A_{F}^{\mathcal{V}_{i}},$
@@
Our goal, then, is to find a way to efficiently and in a generalizable manner, evaluate $A({\mathcal{V}_{i}})$, facilitating Monte Carlo simulations of the coarse-grained system. To achieve this, we use a fully-connected, feed-forward neural network to fit $A({\mathcal{V}_{i}})$ based on the geometry of the Voronoi cell. Therefore, we first have to characterize the Voronoi cell in terms of parameters, which can serve as input for the neural network to predict the free area. Since Voronoi cells can have a varying number of edges, it is challenging to use a fixed number of parameters to capture the cell. To address this, we subdivide each Voronoi cell into triangles, where each triangle connects one of the edges to the center of the cell (see Fig. 3). We can then apply the neural-network fit to each triangle individually, after which the effective potential of the entire Voronoi cell is obtained by summing over the outputs of all triangles. Note that for this system, this is exact, but it would become an approximation for interacting polymers.
@@
Figure 3: a. Cartoon of a Voronoi and Delaunay tessellation for a system of colloids, where the thin, black lines represent the Voronoi tessellation and the thick, green lines represent the Delaunay tessellation. The red dots correspond both to the corners of the Voronoi cells, as well as the centers of the circumscribed circles associated with the Delaunay triangles. b. The subdivision of a Voronoi cell in terms of triangles, where each triangle connects the center of the Voronoi cell with one of the edges. The depletion zone of the Voronoi cell is shown in light green. Each triangle is represented in terms of a set of parameters as shown in the bottom figure. Additionally to the parameters shown in this figure, we also consider the perimeter $P$, the area $A_{T}$ and the ratio ($A_{T}/P^{2}$). Note that the free area $A_{F}$ of the triangle is represented by the white area inside the triangle.
II.3 Fitting the free volume of a Voronoi cell
@@ II.3.1 Training data
    We create a triangle by generating two random numbers between 00 and $1$ that correspond to the sides connected to the particle ($S_{1}$ and $S_{3}$) and a random angle in the range $\theta\in[0,\pi]$ between those two sides. This generates triangles with an essentially arbitrary shape.
@@
As shown in Eq. (10), there is a subset of triangles for which the free area is trivial. In the MC simulation of the coarse grained system, we analytically calculate the free area of those triangles, and only use the neural network to predict the free area for the non-trivial cases. Note, however, that we still included the trivial triangles in our training data because it ensures that the extreme values of the triangles for which the NN has to be used, lie well within the boundaries of the training dataset. As a result, the NN has to extrapolate less at the boundaries. Moreover, in the explicit simulations we saw many triangles with a small $\theta$. Therefore, we ensured that the lower tail of the $\theta$ distribution was well represented in the training data; for every $20,000$ triangles we always included $100$ triangles with a random $\theta<0.0001$ and $100$ triangles with a random $\theta<0.1$. Since we observed that triangles with $\theta\approx\pi$ almost never occur in the actual system, we did not include additional triangles with a large $\theta$ in the training data. To compare the artificially generated training data with the brute-force training data, in Fig. 4 we also show $P(\theta)$ for the artificially generated triangles. As we can see in this figure, the artificially generated triangles lead to a significantly better balanced training data set.
@@ II.4 Implement algorithm in Monte Carlo simulation
In order to be able to test the quality of the ML potential, we have to implement the Voronoi tessellation in an MC simulation. In these MC simulations, trial particle moves are accepted or declined according to the overall Boltzmann weight of the system. For the system under consideration, this Boltzmann weight solely depends on the change in free area of the Voronoi cell of the displaced particle and its neighbors, before and after the move. Simulations of such Voronoi tesselations have been widely used in the context of vertex models to simulate e.g. tissue, and in this paper we build on the developments in this field honda1980much ; bi2016motility ; barton2017active . It is numerically very costly to construct a new Voronoi tessellation (VT) from scratch for each new configurationbarton2017active . However, since the difference between consecutive configurations and thus tessellations in an MC simulation will be minimal, we can instead use the old tessellation as a starting point to obtain the new tessellationbarton2017active . Although we are not aware of an algorithm that directly updates a VT, we can circumvent this problem by turning to the dual of the Voronoi graph, namely the Delaunay tessellation (DT). In the DT, every two particles that share an edge in the VT are connected, such that the DT divides the area into triangles that always connect three particles (see Fig. 3a). Given an arbitrary tessellation, one can converge to the actual DT by applying the equiangulation procedure brakke1992surface . For this, we make use of the fact that in a correct DT the circumscribed circle of each of the triangles cannot contain any other vertices. If two adjacent triangles have opposite angles that are together larger than $\pi$, this circumscribed criterion is no longer met, meaning that the shared edge should flip (see Fig. 5). By iteratively switching edges, it can be shown that the system always converges to the correct DTbrakke1992surface . Given this correct DT, one can then convert back to the VT and compute the free area associated with that configuration. In the SI we discuss in more detail how we realized the implementation of the Delaunay tessellation in an MC simulation.
@@
Figure 6: The pair distribution function $g(r/\sigma_{C})$ plotted for a polymer size of $\sigma_{P}=0.4$ and $\sigma_{P}=0.8$ in respectively panel a) an panel b) for various polymer packing fractions $\eta_{P}\in[0.1,2.5,5.0,8.0]$ and two colloid packing fractions $\eta_{C}\in[0.4,0.6]$. The solid lines represent the ground truth as measured in the system where polymers are treated explicitly whereas the plot markers represent the results as measured in the ML systems. Both the solid lines as well as the plot markers are averaged over five different simulations.
Figure 7: The fraction of free area plotted as a function of the polymer packing fraction $\eta_{P}$ for two polymer sizes, $\sigma_{P}=0.4\sigma_{C}$ (left) and $\sigma_{P}=0.8\sigma_{C}$ (right) and two colloid packing fractions $\eta_{C}=[0.4,0.6]$. The solid lines represent the ground truth as measured in the system where polymers are treated explicitly (here, the small transparent region indicates the standard deviation) The plot markers represent the free area fraction measured in the system simulated with the machine learned potential. For clarity, different radial distributions are shifted vertically.

### v2

@@ Abstract
Machine learning (ML) strategies are opening the door to faster computer simulations, allowing us to simulate more realistic colloidal systems. Since the interactions in colloidal systems are often highly many-body, stemming from e.g. depletion and steric interactions, one of the challenges for these algorithms is capturing the many-body nature of these interactions. In this paper, we introduce a new ML-based strategy for fitting many-body interactions in colloidal systems where the many-body interaction is highly local. To this end, we develop Voronoi-based descriptors for capturing the local environment and fit the effective potential using a simple neural network. To test this algorithm, we consider a simple two-dimensional model for a colloid-polymer mixture, where the colloid-colloid interactions and colloid-polymer interactions are hard-disk like, while the polymers themselves interact as ideal gas particles. We find that a Voronoi-based description is sufficient to accurately capture the many-body nature of this system. Moreover, we find that the Pearson correlation function alone is insufficient to determine the predictive power of the network emphasizing the importance of additional metrics when assessing the quality of ML-based potentials.
@@ I Introduction
Over the last few decades, progress in colloidal particle synthesis has led to an incredible collection of colloidal building blocks ranging from a few nanometers to microns in size, and with an ever-increasing variety of shapes and complex interactions. Together with these impressive strides in colloidal synthesis, there has been an increasing desire to be able to simulate such systems accurately enough to predict their phase behavior, as well as to target their design. However, accurately modeling colloidal systems can be very computationally costly as it requires modeling the colloid core, ligands on their surface, the solvent molecules in which the colloids are dispersed, polymers in the solvent, etc. As such, even accurately simulating a small number of such particles can be computationally expensive. One approach to circumvent these expensive simulations is to integrate out some degrees of freedom associated with the system. This leads to an effective Hamiltonian that is exact, but often includes many-body terms. While such a Hamiltonian can in principle be used to capture the system’s behaviour, the practical evaluation of these higher-order interactions can remain a significant computational bottleneckdijkstra2006effect ; kobayashi2019correction .
One of the possible solutions to speed up the evaluation of these many-body terms – and simulations in general – lies in the realm of machine learning (ML). Over the past decade, ML has emerged as an important tool in speeding up computer simulations of atomic and molecular systems behler2016perspective ; noe2020machine by replacing expensive density functional theory calculations for electronic structure on the atomic scale, with machine learned forces and/or potentials. After training, these ML-based interaction potentials are exploited in either molecular dynamics or Monte Carlo (MC) simulations, allowing researchers to access larger system sizes and longer time scales than would be accessible otherwise. For such systems, strategies are rapidly evolving and improving – from the early “standard” neural network (NN) approaches, to kernel fitting, on the fly training, and message passing neural networks behler2007generalized ; artrith2016implementation ; li2015molecular ; jacobsen2018fly ; jinnouchi2019phase ; kocer2022neural ; batatia2022mace .
@@
We find that for a single size ratio (colloid-to-polymer diameter), a single Voronoi-inspired ML model can reproduce the structure of all relevant phases (gas, liquid, crystal) over all densities. Moreover, we find that the virial pressures in the ML system match the associated brute-force pressures. Additionally, we find that it is difficult to predict the performance of the ML models based on only the correlation between predicted and true effective interactions. In particular, high Pearson correlation coefficients (>0.99999>0.99999) are insufficient to guarantee that the ML-based simulation will accurately reproduce the correct structure of the system. Hence, in this paper, we compare the radial distribution function of the fitted model to results from simulations where the polymers are treated explicitly to assess the quality of the potential.
Figure 2: Probability to have an effective $n$-body colloidal interaction at an arbitrary point in space, for various polymer sizes $\sigma_{P}$ and various colloid and polymer packing fractions $\eta_{C}$ and $\eta_{P}$. Probability is measured using MC simulation of the $N_{C}\mu_{P}AT$ colloid-polymer ensemble, where $N_{C}$ is set to 1024.
@@ II.1 Model system
$\Phi_{\text{eff}}=\frac{1}{2}\sum_{i}^{N_{C}}\sum_{i\neq j}^{N_{C}}\phi_{CC}^{HS}(\left|\mathbf{R}_{i}-\mathbf{R}_{j}\right|)+\Phi_{\text{eff, CP}}(\left\{\mathbf{R}_{i}\right\}),$ (3)
@@
$\Phi_{\text{eff, CP}}(\left\{\mathbf{R}_{i}\right\})$ $=$ $-k_{B}Tz_{P}A_{\text{eff}}(\left\{\mathbf{R}_{i}\right\})$ (4)
$A_{\text{eff}}(\left\{\mathbf{R}_{i}\right\})$ $=$ $\int\mathrm{d}\mathbf{r}\,\exp\left[-\beta\sum_{i=1}^{N_{C}}\phi_{CP}(\left|\mathbf{R}_{i}-\mathbf{r}\right|)\right].$ (5)
@@
Here, $z_{P}$ is the polymer fugacity, $k_{B}$ the Boltzmann constant, $\beta=1/(k_{B}T)$, and $A_{\text{eff}}(\left\{\mathbf{R}_{i}\right\})$ is the effective free area available to the polymers. This effective free area can be interpreted as a weighted area of the system, where the weight of each point in space is determined by the potential energy that a polymer at that position would experience due to the colloid configuration $\left\{\mathbf{R}_{i}\right\}$. Note that Eqs. (3) and (5) are valid for any arbitrary colloid-colloid and colloid-polymer interaction, provided that there are no polymer-polymer interactions.
@@
In the case of hard-sphere colloid-polymer interactions, only points where a polymer can be placed without it overlapping with any of the colloids contribute to $A_{\text{eff, CP}}(\left\{\mathbf{R}_{i}\right\})$. As a result, $A_{\text{eff, CP}}(\left\{\mathbf{R}_{i}\right\})$ simplifies to the free area available to polymers, $A_{F}$.
@@
In this paper, we focus on two polymer sizes: $\sigma_{P}=0.4\sigma_{C}$ and $\sigma_{P}=0.8\sigma_{C}$. These polymer sizes are both far outside of the regime where the approximations of the Asakura-Oosawa pair potential hold, i.e. where the effective colloid-colloid interaction can be written purely as a two-body interaction gast1983polymer ; dijkstra1999phaseb . Hence, for both polymer sizes, $\Phi_{\text{eff, CP}}(\left\{\mathbf{R}_{i}\right\})$ incorporates at least three-body interaction contributions.
@@
Note that the many-body nature of this potential arises due to the fact that more than two depletion zones (disks of radius $\sigma_{CP}/2$ around the colloid where polymer centers cannot enter) can overlap simultaneously, see Fig. 1. The probability $P(n)$ to find $n$ overlapping depletion zones at an arbitrary point in space depends on the colloid and polymer concentration, as well as the polymer size, as is shown in Fig. 2. The data in this figure is obtained by performing an MC simulation in the $N_{C}\mu_{P}AT$ ensemble, where polymers were treated explicitly. To ensure a constant $\mu_{P}$, we perform insertion and deletion moves of polymers via the usual Monte Carlo acceptance rule frenkel2023understanding . To compute $P(n)$, we randomly selected points within the system and determined the number of depletion zones overlapping with these points (where points located within the colloids were included in the computation as well). For each state point, we analyzed 45 different snapshots, where we considered $10^{6}$ random points per snapshot. From Fig. 2 we see that at lower polymer sizes there are mainly $1$- and $2$-body interactions, while at higher polymer sizes the system mainly shows $3$-body interactions.
@@ II.2 Capture effective potential using Voronoi cells
Our aim is to train an ML algorithm that is able to predict the effective interactions between colloids as described in Eq. (5). However, rather than predicting the free area for the entire system at once, we aim to compute the contributions to the effective potential of individual colloids. This requires rewriting $A_{\text{eff}}(\left\{\mathbf{R}_{i}\right\})$ in terms of contributions associated with separate colloids, i.e. $A_{\text{eff}}(\left\{\mathbf{R}_{i}\right\})=\sum_{i\in N_{C}}A^{i}_{\text{eff}}(\left\{\mathbf{R}_{i}\right\})$. To achieve this, each point in space must be assigned to a specific colloid, allowing the integral of Eq. (5) to be split into $N_{C}$ integrals. A natural way to tile space is by using Voronoi cells, where each point in space is assigned to the nearest colloid, allowing us to write
$A_{\text{eff}}(\left\{\mathbf{R}_{i}\right\})=\sum_{i\in N_{C}}\int_{\mathcal{V}_{i}}\mathrm{d}\mathbf{r}\,\exp\left[-\beta\sum_{j}^{N_{C}}\phi_{CP}(\mathbf{R}_{j}-\mathbf{r})\right],$
@@
$A_{\text{eff}}(\left\{\mathbf{R}_{i}\right\})=\sum_{i\in N_{C}}A^{(1)}({\mathcal{V}_{i}})+\sum_{i,j\in N_{C}}A^{(2)}({\mathcal{V}_{i}},{\mathcal{V}_{j}})+\ldots,$ (6)
@@
$A_{\text{eff}}(\left\{\mathbf{R}_{i}\right\})=\sum_{i\in N_{C}}A_{F}^{\mathcal{V}_{i}},$
@@
Our goal, then, is to find a way to efficiently and in a generalizable manner, evaluate $A({\mathcal{V}_{i}})$, facilitating Monte Carlo simulations of the coarse-grained system. To achieve this, we use a fully-connected, feed-forward neural network to fit $A({\mathcal{V}_{i}})$ based on the geometry of the Voronoi cell. Therefore, we first have to characterize the Voronoi cell in terms of parameters, which can serve as input for the neural network to predict the free area. Since Voronoi cells can have a varying number of edges, it is challenging to use a fixed number of parameters to capture the cell. To address this, we subdivide each Voronoi cell into triangles, where each triangle connects one of the edges to the center of the cell (see Fig. 3). We can then apply the neural-network fit to each triangle individually, after which the effective potential of the entire Voronoi cell is obtained by summing over the outputs of all triangles. Note that for this system, this approach is exact, but it would become an approximation for interacting polymers.
@@
Figure 3: a. Cartoon of a Voronoi and Delaunay tessellation for a system of colloids, where the thin, black lines represent the Voronoi tessellation and the thick, green lines represent the Delaunay tessellation. The red dots correspond both to the corners of the Voronoi cells, as well as the centers of the circumscribed circles associated with the Delaunay triangles. b. The subdivision of a Voronoi cell in terms of triangles, where each triangle connects the center of the Voronoi cell with one of the edges. The depletion zone of the Voronoi cell is shown in light green. Each triangle is represented in terms of a set of parameters as shown in the bottom figure. Additionally to the parameters shown in this figure, we also consider the perimeter $P$, the area $A_{T}$ and the ratio ($A_{T}/P^{2}$). Note that the free area $A_{F}$ of the triangle is equal to the white area inside the triangle.
II.3 Fitting the free area of a Voronoi cell
@@ II.3.1 Training data
    We create a triangle by generating two random numbers between 0 and $1$ that correspond to the sides connected to the particle ($S_{1}$ and $S_{3}$) and a random angle in the range $\theta\in[0,\pi]$ between those two sides. This generates triangles with an essentially arbitrary shape.
@@
As shown in Eq. (10), there is a subset of triangles for which the free area is trivial. In the MC simulation of the coarse grained system, we analytically calculate the free area of those triangles, and thus only use the neural network to predict the free area for the non-trivial cases. Note, however, that we still included the trivial triangles in our training data because it ensures that the extreme values of the triangles for which the NN has to be used, lie well within the boundaries of the training dataset. As a result, the NN has to extrapolate less at the boundaries. Moreover, in the explicit simulations we saw many triangles with a small $\theta$. Therefore, we ensured that the lower tail of the $\theta$ distribution was well represented in the training data; for every $20,000$ triangles we always included $100$ triangles with a random $\theta<0.0001$ and $100$ triangles with a random $\theta<0.1$. Since we observed that triangles with $\theta\approx\pi$ almost never occur in the actual system, we did not include additional triangles with a large $\theta$ in the training data. To compare the artificially generated training data with the brute-force training data, in Fig. 4 we also show $P(\theta)$ for the artificially generated triangles. As we can see in this figure, the artificially generated triangles lead to a significantly better balanced training data set.
@@ II.4 Implement algorithm in Monte Carlo simulation
In order to be able to test the quality of the ML potential, we have to implement the Voronoi tessellation in an MC simulation. In these MC simulations, trial particle moves are accepted or declined according to the overall Boltzmann weight of the system. For the system under consideration, this Boltzmann weight solely depends on the change in free area of the Voronoi cell of the displaced particle and its neighbors, before and after the move. Simulations of Voronoi tesselations have been widely used in the context of vertex models to simulate e.g. tissue, and in this paper we build on the developments in this field honda1980much ; bi2016motility ; barton2017active . It is numerically very costly to construct a new Voronoi tessellation (VT) from scratch for each new configurationbarton2017active . However, since the difference between consecutive configurations and thus tessellations in an MC simulation will be minimal, we can instead use the old tessellation as a starting point to obtain the new tessellationbarton2017active . Although we are not aware of an algorithm that directly updates a VT, we can circumvent this problem by turning to the dual of the Voronoi graph, namely the Delaunay tessellation (DT). In the DT, every two particles that share an edge in the VT are connected, such that the DT divides the area into triangles that always connect three particles (see Fig. 3a). Given an arbitrary tessellation, one can converge to the actual DT by applying the equiangulation procedure brakke1992surface . This procedure makes use of the fact that in a correct DT the circumscribed circle of each of the triangles cannot contain any other vertices. If two adjacent triangles have opposite angles that are together larger than $\pi$, this circumscribed criterion is no longer met, meaning that the shared edge should flip (see Fig. 5). By iteratively switching edges, it can be shown that the system always converges to the correct DTbrakke1992surface . Given this correct DT, one can then convert back to the VT and compute the free area associated with that configuration. In the SI we discuss in more detail how we realized the implementation of the Delaunay tessellation in an MC simulation.
@@
Figure 6: The pair distribution function $g(r/\sigma_{C})$ plotted for a polymer size of $\sigma_{P}=0.4$ and $\sigma_{P}=0.8$ in respectively panel a) and panel b) for various polymer packing fractions $\eta_{P}\in[0.1,2.5,5.0,8.0]$ and two colloid packing fractions $\eta_{C}\in[0.4,0.6]$. The solid lines represent the ground truth as measured in the system where polymers are treated explicitly whereas the plot markers represent the results as measured in the ML systems. Both the solid lines as well as the plot markers are averaged over five different simulations. For clarity, the radial distribution functions are shifted vertically, with higher curves corresponding to systems with larger polymer fractions.
Figure 7: The fraction of free area plotted as a function of the polymer packing fraction $\eta_{P}$ for two polymer sizes, $\sigma_{P}=0.4\sigma_{C}$ (left) and $\sigma_{P}=0.8\sigma_{C}$ (right) and two colloid packing fractions $\eta_{C}=[0.4,0.6]$. The lines represent the ground truth as measured in the system where polymers are treated explicitly (here, the small transparent region indicates the standard deviation) The plot markers represent the free area fraction measured in the system simulated with the ML potential.
Figure 8: The pressure plotted as a function of the polymer packing fraction $\eta_{P}$ for two polymer sizes, $\sigma_{P}=0.4\sigma_{C}$ (left) and $\sigma_{P}=0.8\sigma_{C}$ (right) and two colloid packing fractions $\eta_{C}=[0.4,0.6]$. The lines represent the ground truth as measured in the system where polymers are treated explicitly (here, the small transparent region indicates the standard deviation). The plot markers represent the pressure measured in the system simulated with the machine learned potential.